OpenStudy (anonymous):

Help?!?!? What is the solution of the following system? Show your work. 3x + 3y = 10 -9x - 9y = -30

OpenStudy (anonymous):

Sorry about the typo. Its supposed to look like this. 3x + 3y = 10 -9x - 9y = -30

OpenStudy (anonymous):

\[3x+3y=10\] \[-3x-3y=-10...(2)\]

OpenStudy (cwrw238):

3x + 3y = 10 -9x - 9y = -30 multiply first equation by 3 9x + 9y = 30 adding gives 0 = 0

OpenStudy (cwrw238):

yes this means infinite solutions

OpenStudy (anonymous):

Oh i see... thanks! @cwrw238

OpenStudy (anonymous):

Are you sure thats the answer? Cause 0 = 0 doesnt look right...

OpenStudy (anonymous):

@cwrw238